# Self-Rewarding Language Models

### v1

@@ 3.2.1 Instruction Following Ability
3.2.2 Reward Modeling Ability
Reward modeling evaluation results are provided in Table 2.
EFT augmentation improves over SFT baseline
Firstly, we find that adding Evaluation Fine-Tuning (EFT) data into training, which gives examples to the model of how to act as an LLM-as-a-Judge, naturally improves its performance compared to training with Instruction Fine-Tuning (IFT) data alone. IFT data covers a wide range of general instruction tasks, and so does endow the SFT Baseline with the ability to evaluate responses, however EFT data gives more examples of this specific task. We find improvements across all five metrics measured when using IFT+EFT vs. IFT alone, e.g. the pairwise accuracy agreement with humans increases from 65.1% to 78.7%.
@@ 3.2.2 Reward Modeling Ability
In these experiments we used the LLM-as-Judge prompt format shown in Figure 2. In preliminary experiments we also tried various other prompts to decide the most effective one to use. For example, we tried the prompt proposed in Li et al. (2023a) which also proposes a 5-point scale, but describes the options as multiple choice in a range of quality buckets, see Figure 5. In contrast, our prompt describes the points as additive, covering various aspects of quality. We find a large difference between these two prompts when using the SFT Baseline, e.g. 65.1% pairwise accuracy for ours, and only 26.6% pairwise accuracy for theirs. See Section A.2 for further details.
@@ 4 Related Work
Preference learning approaches such as in Ziegler et al. (2019); Stiennon et al. (2020); Ouyang et al. (2022); Bai et al. (2022a) train a fixed reward model from human preference data, and then use the reward model to train via reinforcement learning (RL), e.g. via Proximal Policy Optimization (PPO) (Schulman et al., 2017). Thus, the reward signal in a certain sense already comes from a model even in these works, but distilled from human data. Nevertheless, this is commonly referred to as RL from Human Feedback (RLHF). Methods such as Direct Preference Optimization (DPO) (Rafailov et al., 2023) avoid training the reward model entirely, and instead directly train the LLM using human preferences. Several other such competing methods exist as well (Zhao et al., 2023; Zheng et al., 2023a; Yuan et al., 2023), including Pairwise Cringe Optimization (PCO) (Xu et al., 2023), which was shown to outperform DPO and PPO on AlpacaFarm (Dubois et al., 2023). PCO uses an iterative training approach similar to the one in our work, except with a fixed reward model, and also showed that Iterative DPO improves over DPO using the same scheme.
@@
Constitutional AI (Bai et al., 2022b) uses an LLM to give feedback and refine responses, and uses this data to train a (separate, fixed) reward model. This is then used to perform “RL from AI Feedback” (RLAIF). Lee et al. (2023) compare RLAIF and RLHF procedures and find the methods they compare perform roughly equally. They use an “off-the-shelf” LLM to perform LLM-as-a-Judge prompting to build a training set to train a fixed reward model, which is then used for RL training. They also experiment with using the fixed but separate LLM-as-a-Judge model directly, which the authors report is computationally expensive due to using it within PPO training (rather than the offline step in the iterative approach we use in our work, which is relatively computationally cheap). Finally, Chen et al. (2024) recently showed they can avoid reward models entirely in an Iterative DPO-like framework by using human labels as the winning response in a pair, and the last iteration’s generations as the losing response in the pair. The authors note this has the limitation that once the model generations reach human performance, they are bottlenecked. Further, each input prompt is required to have a human annotated response, in contrast to our work.
@@
Several methods have improved LLMs by (self-)creating training data to augment fine-tuning. Self-Instruct (Wang et al., 2022) is a method for self-instruction creation of prompts and responses, which can be used to improve a base LLM. We make use of a similar technique in our work, and then use our self-reward model to score them. Several approaches have also created training data by distilling from powerful LLMs, and shown a weaker LLM can then perform well. For example, Alpaca (Taori et al., 2023) fine-tuned a Llama 7B model with text-davinci-003 instructions created in the style of self-instruct. Alpagasus (Chen et al., 2023) employed a strong LLM-as-a-Judge (ChatGPT) to curate the Alpaca dataset and filter to a smaller set, obtaining improved results. Instruction Backtranslation (Li et al., 2023a) similarly augments and curates training data, but augmenting via backtranslating from web documents to predict prompts. The curation is done by the LLM(-as-a-Judge) itself, so can be seen as an instance of a self-rewarding model, but in a specialized setting. Reinforced Self-Training (ReST) (Gulcehre et al., 2023) uses a fixed, external reward to curate new high-quality examples to iteratively add to the training set, improving performance. We note that in our experiments, we found that adding only positive examples in a related manner did not help, whereas adding preference pairs did help.
@@
Using LLM-as-a-Judge prompting to evaluate language models has become a standard approach (Dubois et al., 2023; Li et al., 2023b; Fernandes et al., 2023; Bai et al., 2023; Saha et al., 2023), and is being used to train reward models or curate data as well, as described above (Lee et al., 2023; Chen et al., 2023; Li et al., 2023a). While some works such as Kim et al. (2023) create training to data to train an LLM to perform well as a judge, to our knowledge it is not common to combine this training with general instruction following skills as in our work.
@@ 5 Conclusion
We have introduced Self-Rewarding Language Models, models capable of self-alignment via judging and training on their own generations. The method is trained in an iterative manner, where in each iteration the model creates its own preference-based instruction training data. This is done by assigning rewards to its own generations via LLM-as-a-Judge prompting, and using Iterative DPO to train on the preferences. We showed that this training both improves the instruction following capability of the model, as well as its reward-modeling ability across the iterations. While this is only a preliminary study, we believe this is an exciting avenue of research because this means the model is better able to assign rewards in future iterations for improving instruction following – a kind of virtuous circle. While this improvement likely saturates in realistic scenarios, it still allows for the possibility of continual improvement beyond the human preferences that are typically used to build reward models and instruction following models today.
@@ 6 Limitations
While we have evaluated our models using GPT-4 using head-to-head and AlpacaEval 2 leaderboard style evaluation, there are many other automatic evaluation benchmarks that one can measure. Further, we observed an increase in length in model generations, and there is a known correlation between length and estimated quality, which is a topic that should be understood more deeply in general, and in our results in particular as well. It would also be good to understand if so-called “’reward-hacking” can happen within our framework, and in what circumstances. As we are using both a language model as the training reward, and a language model for final evaluation, even if they are different models, this may require a deeper analysis than we have provided. We conducted a preliminary human (author) evaluation, which validated the automatic results that we see, but more detailed human evaluation would be beneficial.
Another clear further avenue of study is to conduct safety evaluations – and to explore safety training within our framework. Reward models have been built exclusively for safety in existing systems (Touvron et al., 2023), and a promising avenue here would be to use the LLM-as-a-Judge procedure to evaluate for safety specifically in our self-rewarding training process. Given that we have shown that reward modeling ability improves over training iterations, this could mean that the safety of the model could potentially improve over time as well, with later iterations being able to capture more challenging safety situations that earlier iterations cannot.
@@ A.1 Distributions of IFT, EFT and AIFT data
We have plotted the distribution of instructions for IFT, EFT and AIFT data, and the distribution of responses for IFT, EFT and AIFT data in Figure 4. It is clear that the IFT data and EFT data come from very different distributions while the IFT and AIFT data come from similar distributions.
A.2 Other EFT prompts we have tried
At first, we took the EFT prompt from Li et al. [2023a] as shown in Figure 5. However, we found that this prompt was not as effective as our additive score-counting prompt because the model needed to treat the task as a multiple-choice problem, and it was difficult for the model to break down this multiple-choice problem into sub-problems involving evaluating various aspects of the response. When using the model trained on 3,200 IFT data only, its performance on the EFT test set using our additive score-counting prompt and prompt from Li et al. [2023a] is shown in Table 3.
@@ A.2 Other EFT prompts we have tried
<response><RESPONSE_HERE></response>
@@
Figure 5: LLM-as-a-Judge prompt taken from Li et al. [2023a].
@@
Table 3: We tried various LLM-as-Judge prompts using the model trained with 3,200 IFT data only and found that our additive score-counting prompt worked best which demonstrates significant improvements in EFT performance comparing to the prompt used by Li et al. [2023a].

### v3

@@ 3.2.2 Reward Modeling Ability
In these experiments we used the LLM-as-a-Judge prompt format shown in Figure 2. In preliminary experiments we also tried various other prompts to decide the most effective one to use. For example, we tried the prompt proposed in Li et al. (2024) which also proposes a 5-point scale, but describes the options as multiple choice in a range of quality buckets, see Appendix Figure 7. In contrast, our prompt describes the points as additive, covering various aspects of quality. We find a large difference between these two prompts when using the SFT Baseline, e.g. 65.1% pairwise accuracy for ours, and only 26.6% pairwise accuracy for theirs. See Appendix A.2 for further details.
@@ 4 Related Work
Preference learning approaches such as in Ziegler et al. (2019); Stiennon et al. (2020); Ouyang et al. (2022); Bai et al. (2022a) train a fixed reward model from human preference data, and then use the reward model to train via reinforcement learning (RL), e.g. via Proximal Policy Optimization (PPO) (Schulman et al., 2017). Thus, the reward signal in a certain sense already comes from a model even in these works, but distilled from human data. Nevertheless, this is commonly referred to as RL from Human Feedback (RLHF). Methods such as Direct Preference Optimization (DPO) (Rafailov et al., 2023) avoid training the reward model entirely, and instead directly train the LLM using human preferences. Several other such competing methods exist as well (Zhao et al., 2023; Zheng et al., 2023a; Yuan et al., 2023), including Pairwise Cringe Optimization (PCO) (Xu et al., 2023). PCO uses an iterative training approach similar to the one in our work, except with a fixed reward model, and that work also showed that Iterative DPO improves over DPO using the same scheme. We note that other works have developed iterative preference training schemes as well, e.g. Adolphs et al. (2023); Gulcehre et al. (2023); Xiong et al. (2023).
@@
Constitutional AI (Bai et al., 2022b) uses an LLM to give feedback and refine responses, and uses this data to train a reward model. This fixed, separate reward model is then used to train the language model via RL, called “RL from AI Feedback” (RLAIF). Lee et al. (2023) compare RLAIF and RLHF procedures and find the methods they compare perform roughly equally. They use an “off-the-shelf” LLM to perform LLM-as-a-Judge prompting to build a training set to train a fixed reward model, which is then used for RL training. They also experiment with using the fixed but separate LLM-as-a-Judge model directly, which the authors report is computationally expensive due to using it within PPO training (rather than the offline step in the iterative approach we use in our work, which is relatively computationally cheap). Finally, SPIN (Chen et al., 2024b) recently showed they can avoid reward models entirely in an Iterative DPO-like framework by using human labels as the winning response in a pair, and the last iteration’s generations as the losing response in the pair. The authors note this has the limitation that once the model generations reach human performance, they are bottlenecked. Further, each input prompt is required to have a human annotated response, in contrast to our work.
@@
Several methods have improved LLMs by (self-)creating training data to augment fine-tuning. Self-Instruct (Wang et al., 2023) is a method for self-instruction creation of prompts and responses, which can be used to improve a base LLM. We make use of a similar technique in our work, and then use our self-reward model to score them. Several approaches have also created training data by distilling from powerful LLMs, and shown a weaker LLM can then perform well. For example, Alpaca (Taori et al., 2023) fine-tuned a Llama 7B model with text-davinci-003 instructions created in the style of self-instruct. Alpagasus (Chen et al., 2024a) employed a strong LLM-as-a-Judge (ChatGPT) to curate the Alpaca dataset and filter to a smaller set, obtaining improved results. Instruction Backtranslation (Li et al., 2024) similarly augments and curates training data, but augmenting via backtranslating from web documents to predict prompts. The curation is done by the LLM(-as-a-Judge) itself, so can be seen as an instance of a self-rewarding model, but in a specialized setting. Reinforced Self-Training (ReST) (Gulcehre et al., 2023) uses a fixed, external reward to curate new high-quality examples to iteratively add to the training set, improving performance. In our experiments, we found that adding only positive examples in a related manner did not help, whereas preference pairs did help (see Appendix Section A.4 for details).
@@
Using LLM-as-a-Judge prompting to evaluate language models has become a standard approach (Dubois et al., 2023; Li et al., 2023; Fernandes et al., 2023; Bai et al., 2023; Saha et al., 2023), and is being used to train reward models or curate data as well, as described above (Lee et al., 2023; Chen et al., 2024a; Li et al., 2024). While some works such as Kim et al. (2023) create training data to train an LLM to perform well as a judge, to our knowledge it is not common to combine this training with general instruction following skills as in our work.
@@ 5 Conclusion
We have introduced Self-Rewarding Language Models, models capable of self-alignment via judging and training on their own generations. The method learns in an iterative manner, where in each iteration the model creates its own preference-based instruction training data. This is done by assigning rewards to its own generations via LLM-as-a-Judge prompting, and using Iterative DPO to train on the preferences. We showed that this training both improves the instruction following capability of the model, as well as its reward-modeling ability across the iterations. While there are many avenues left unexplored, we believe this is exciting because this means the model is better able to assign rewards in future iterations for improving instruction following – a kind of virtuous circle. While this improvement likely saturates in realistic scenarios, it still allows for the possibility of continual improvement beyond the human preferences that are typically used to build reward models and instruction following models today.
@@ 6 Limitations
We observed an increase in length in model generations, and there is a known correlation between length and estimated quality, which is a topic that should be understood more deeply in general, and in our results in particular as well. It would also be good to understand if so-called “reward-hacking” can happen within our framework, and in what circumstances. As we are using both a language model as the training reward, and a language model for final evaluation (GPT-4) in some of our benchmarks, even if they are different models, this may require a deeper analysis than we have provided. While the human evaluation we conducted did provide validation of the automatic results, further study could bring more insights.
Another clear further avenue of study is to conduct safety evaluations – and to explore safety training within our framework. Reward models have been built exclusively for safety in existing systems (Touvron et al., 2023), and a promising avenue here would be to use the LLM-as-a-Judge procedure to evaluate for safety specifically in our self-rewarding training process. Given that we have shown that reward modeling ability improves over training iterations, this could mean that the safety of the model could potentially improve over time as well, with later iterations being able to catch and mitigate more challenging safety situations that earlier iterations cannot.
@@ A.1 Distributions of IFT, EFT and AIFT data
We have plotted the distribution of instructions for IFT, EFT and AIFT($M_{1}$) data, and the distribution of responses for IFT, EFT and AIFT($M_{1}$) data in Figure 6. It is clear that the IFT data and EFT data come from very different distributions while the IFT and AIFT($M_{1}$) data come from similar distributions.
A.2 EFT Prompts
The EFT prompt which we use in our main experiments is shown in Figure 2.
Other EFT prompts we have tried
At first, we took the EFT prompt from Li et al. [2024] as shown in Figure 7. However, we found that this prompt was not as effective as our additive score-counting prompt because the model needed to treat the task as a multiple-choice problem, and it was difficult for the model to break down this multiple-choice problem into sub-problems involving evaluating various aspects of the response. When using the model trained on 3,200 IFT data only, its performance on the EFT test set using our additive score-counting prompt and prompt from Li et al. [2024] is shown in Table 5.
@@ A.2 EFT Prompts
response><RESPONSE_HERE></response>
@@
Figure 7: LLM-as-a-Judge prompt taken from Li et al. [2024].
@@
Table 5: We tried various LLM-as-Judge prompts using the model trained with 3,200 IFT data only and found that our additive score-counting prompt worked best which demonstrates significant improvements in EFT performance comparing to the prompt used by Li et al. [2024].
A.3 Self-rewarding Models Using IFT Data Only
To demonstrate the importance of the EFT data, we also trained a series of models starting with the model trained only on the IFT data. The following is the model sequence.
  • $M_{0}$
    : Base pretrained LLM with no fine-tuning.
  • $M_{1}^{\prime}$
    : Initialized with $M_{0}$, then fine-tuned on the IFT seed data only using SFT.
  • $M_{2}^{\prime}$
    : Initialized with $M_{1}^{\prime}$, then trained with AIFT($M_{1}^{\prime}$) data using DPO.
  • $M_{3}^{\prime}$
    : Initialized with $M_{2}^{\prime}$, then trained with AIFT($M_{2}^{\prime}$) data using DPO.
Since we did not use EFT data to train the series of models, they were not always able to score the responses according to the format and even when they did, the scores given typically converged to 4. Therefore, even when starting from the same number of generated new prompts, we could only collect a very small number of valid training samples for DPO. In total, we collected 541 pairs to form the AIFT($M_{1}^{\prime}$) dataset used to train $M_{2}^{\prime}$ via DPO, and 429 pairs to form AIFT($M_{2}^{\prime}$) used to train $M_{3}^{\prime}$. The win rates are shown in Figure 8. From the figure we can conclude that EFT data helps to get better performance in the same number of iterations and the gap in performance between the model trained with EFT data and the model trained without EFT data widens in the later iterations.
Figure 8: EFT data helps the self-rewarding loop: We evaluated the series of models trained using self-reward loops starting from the model trained using only IFT data. We performed head-to-head win rates comparisons on the IFT test set. While $M_{2}^{\prime}$ can improve over the SFT baseline and $M_{3}^{\prime}$ can improve even more over the SFT baseline, they lag far behind the corresponding models ($M_{2}$, $M_{3}$) that started from a base model trained using both IFT and EFT data, see Figure 3.
<LIST ALL ALPACAEVAL INSTRUCTIONS>
Given the above list of possible instructions, define a maximum of 20 categories that would cover the types of intructions, for example recipes, reasoning tasks, general knowledge etc. Try to cover as many of the instructions as possible with the maximum 20 categories, while keeping the categories high-level, simple and easy to understand.
Figure 9: Prompt used to obtain instruction categories on the AlpacaEval test set.
Instruction: <INSTRUCTION> Given the above, categorize it into one of the following 20 categories: <LIST ALL CATEGORIES>
Secondly, score the instruction in terms of complexity: how complex you think it is to answer from 1-10 (where 10 is a complex question whereby first reasoning or breaking down the question into multiple subquestions for example might help improve the answer).
Thirdly, indicate how long you think the response to the instruction should be, either (a) 1 sentence, (b) 1-3 sentences, (c) 1 paragraph, (d) 2 paragraphs, or (e) 3 or more paragraphs. Provide your final response in the following format: Category: <one of the 20 categories>
Complexity: <score out of 10>
Length: <length category>. Do not provide the actual response.
Figure 10: Prompt for categorizing instructions based on their topics, complexities and expected response lengths.
Table 6: Breakdown of AlpacaEval test set instructions by instruction category.
Category Number Percentage
Science / Technology / Engineering 134 16.65%
Professional / Business / Marketing 77 9.57%
Social Interaction / Relationships / Human Behavior 68 8.45%
Miscellaneous / Other 61 7.58%
Mathematics / Logical Reasoning 52 6.46%
Cooking / Recipes 48 5.96%
Software Development / Coding / Algorithms 44 5.47%
Travel / Geography / Exploration 41 5.09%
Literature / Writing / Communication 39 4.84%
History / Social Studies 38 4.72%
Entertainment / Media Analysis 34 4.22%
Language Learning / Linguistics 32 3.98%
Music / Audio / Arts 30 3.73%
DIY Projects / Hobbies 24 2.98%
Technology / Gadgets / Consumer Products 20 2.48%
Gaming / Game Development 18 2.24%
Exercise / Health / Wellness 16 1.99%
Philosophy / Ethics / Ideology 15 1.86%
Sports / Athletics / Physical Activity 12 1.49%
Strategy / Problem-Solving / Critical Thinking 2 0.24%
Table 7: Breakdown of AlpacaEval test set instructions by instruction complexity. The instructions increase in complexity from 1 to 9, where 10 is a complex question that requires first reasoning or breaking the problem into sub-problems before it can be solved.
Complexity Number Percentage
3 238 29.57%
2 206 25.59%
4 122 15.16%
6 79 9.81%
5 68 8.45%
7 41 5.09%
1 34 4.22%
8 14 1.74%
9 3 0.37%
Table 8: Breakdown of AlpacaEval test set instructions by expected response length.
Expected Length Number Percentage
1-3 sentences 361 44.84%
1 paragraph 269 33.42%
1 sentence 143 17.76%
2 paragraphs 31 3.85%
3 or more paragraphs 1 0.13%
Figure 11: AlpacaEval win rate breakdown for instruction complexities (left) and expected response lengths (right). Self-Rewarding models give gains across most complexities and all response length ranges.
A.4 Preference optimization outperforms augmenting with positive examples only
We also tried an alternative self-training procedure of adding high-quality self-instruction creation examples to supervised fine-tuning (without preference optimization), rather than DPO. In this variant, we add additional examples of (instruction prompt, response) curated by the model to the seed set for supervised fine-tuning, following other approaches [Li et al., 2024, Adolphs et al., 2023, Gulcehre et al., 2023], rather than constructing preference data. In this setup we only add examples where the candidate response was evaluated to give a perfect score of $r_{i}^{n}=5$. Unfortunately we could not find a configuration where this approach helped. For example, adding 11,254 such examples that scored 5 out of 5, and optimizing the mixing weight in training, still yielded a head to head with the SFT Baseline of 29% wins vs 30% wins, i.e., no improvement.
A.5 Augmented Prompt Generation Using Newly Trained Models
In our experiments, for time efficiency, we have created a fixed pool of augmented prompts in advance using ChatLlama 70B. In a real interactive system, ideally, those prompts could come from real users so that we can ensure the models are trained to align with real user requirements. Here, we also examine whether our newly trained Self-Rewarding models in each iteration can generate new prompts through in-context learning, instead of using ChatLlama 70B. To check this, we constructed 30 prompts with in-context examples using the original seed IFT data as described in Section 2.2 and tested whether $M_{1}$, $M_{2}$ and $M_{3}$ still possess in-context learning ability and can generate high quality instructions. According to manual inspection, all models can generate novel instructions given in-context examples in all 30 cases. However, for $M2$ and $M3$, the model is likely to first generate a few instructions, then generate a separator, and then start responding to the instructions, so some postprocessing might be necessary.
A.6 AlpacaEval Test Sample Clustering
We used the GPT-4 (gpt-4-1106-preview) model to categorize the instructions in the AlpacaEval test set into clusters from three perspectives: (1) instruction category, (2) instruction complexity, and (3) expected response length. To obtain instruction categories for the AlpaceEval test set, we used the prompt in Figure 9 and obtained 20 categories in total. Then, to cluster the instructions into different groups, we use the prompt in Figure 10 for each test example. The corresponding statistics are given in Table 6, Table 7, Table 8. The fine-grained results on instruction complexity and expected response length are given in Figure 11.
Table 9: NLP Benchmarks. Self-Rewarding models mostly tend to maintain performance compared to the Llama 2 base model and the SFT Baseline, despite being fine-tuned on very different instruction-following prompts.
Commonsense Reasoning Math World Knowledge
ARC
easy
ARC
challenge
HellaSwag SIQA PIQA
GSM8K
MMLU
(macro_avg/acc)
OBQA
(acc_comp)
NQ
Llama 2 80.20 57.40 85.30 50.70 82.80 56.80 68.90 60.20 25.30
SFT Baseline 76.49 55.97 85.17 51.48 82.59 50.72 69.76 57.80 34.35
$M_{1}$ 78.14 57.51 84.99 53.02 82.92 60.27 69.34 57.60 35.48
$M_{2}$ 74.84 54.51 84.27 51.23 81.94 59.29 69.31 57.60 33.07
$M_{3}$ 72.35 53.13 83.29 49.28 80.79 57.70 69.37 58.40 31.86
Table 10: MT-Bench Fine-grained Results. We list our models’ performance on each problem category. Self-reward is especially effective in improving the model’s ability in writing, role-playing, extraction, and STEM tasks.
Writing Roleplay Reasoning Math Coding Extraction STEM Humanities Overall
SFT 8.83 8.15 5.30 3.00 3.50 6.90 9.18 9.95 6.85
M1 9.10 7.65 4.35 3.05 4.10 7.20 8.93 9.85 6.78
M2 9.10 8.00 4.60 3.30 4.25 7.65 9.40 9.80 7.01
M3 9.58 8.73 4.80 3.50 4.20 7.80 9.45 9.95 7.25
A.7 NLP Benchmark Results and MT-Bench Results
We provide the detailed model performance on a number of NLP benchmarks in Table 9 and on MT-Bench in Table 10. In particular, some NLP benchmarks including ARC-Challenge, HellaSwag, SIQA, PIQA, and OBQA are all text completion tasks. In these tasks, given the multiple choice options, we choose the option corresponding to the highest log probability scored by the models as the final answer. As such, the objective of these particular tasks is quite different from what our algorithm tries to optimize, so the results on these tasks may not reflect the true capability of our models.